This is an open ended question but I am not sure how to even start this.
The effect of depreciation can be computed using a formula similar to the formula for compound interest. A) assume depreciation is the same each mont. Write a problem involving depreciation and solve it.
B) Develop a general formula for depreciation defining what each variable in the formula stands for.

Answer by stanbon(57214)

The effect of depreciation can be computed using a formula similar to the formula for compound interest.
----------------------------------------
Appreciation: A(t) = P(1+(r/n))^((nt)
---
A) assume depreciation is the same each month. Write a problem involving depreciation and solve it.
----------------------------------
A person inherits 10,000 and decides to spend 6% each quarter of the year.
How much will be left in 5 years?
Depreciation: A(t) = P(1-(r/n))^(nt)
A(5) = 10,000(1-(0.06/4))^(4*5)
A(5) = 10,000(1-0.0150)^20
A(5) = 10,000(0.9850)^20
A(5) = 10,000*0.7391
A(5) = $7391.00
=================================

B) Develop a general formula for depreciation defining what each variable in the formula stands for.
Comment: I'll leave that to you.
======================================
Cheers,
Stan H.